# Statistical Finite Elements via Interacting Particle Langevin Dynamics

### v1

@@ 6 Experimental Results
In this section, we provide comprehensive numerical evidence that the methods we presented above are effective in practice. We consider three examples, each with a different level of complexity, to demonstrate the effectiveness of the IPLA algorithm within the setting of inverse problems. Specifically, we consider the following examples:
@@
  • Nonlinear Poisson forcing estimation: This example introduces a nonlinearity into the Poisson diffusivity, where we can compare different nonlinear approximation methods. Here, we compare the first-order Taylor, unscented transform and Monte Carlo approximations for approximating nonlinear transforms of Gaussian random variables.
  • Linear Poisson inverse problem: Finally, we consider the estimation of an unknown constant diffusivity for the linear Poisson problem, demonstrating the effect of a warm-start on the efficiency of estimating the parameter.
@@ 6.1 Forcing estimation for Linear Poisson PDE
Here we consider two example domains, the unit interval, and a 2D disc. In both cases we fix the diffusivity to a constant value of one. For the unit interval the linear Poisson PDE is given by
@@
with $f(\mathbf{x})=100\exp\left(-\frac{\|\mathbf{x}+\mathbf{m}\|^{2}_{2}}{2*0.2^{2}}\right)+100\exp\left(-\frac{\|\mathbf{x}-\mathbf{m}\|^{2}_{2}}{2*0.2^{2}}\right)$, $\mathbf{m}=[0.4,0.4]^{\top}$, and $k(\mathbf{x},\mathbf{x}^{\prime})=\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}_{2}}{2*0.1^{2}}\right)$. For inference $f\sim\mathcal{GP}(0,k_{f})$, $k_{f}(\mathbf{x},\mathbf{x}^{\prime})=100\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}_{2}}{2*0.2^{2}}\right)$.
@@
The convergence of the estimated forcing to the MMAP estimator is computed by assuming a convergence of the form $\|\mathbf{b}^{(N)}_{k}-\mathbf{b}^{\star}\|_{2}=CN^{-p}$, where $\mathbf{b}^{(N)}_{k}$ is the parameter estimate after $k$ iterations of preconditioned IPLA (27) using $N$ particles. The rate of convergence, $C$ will depend on the problem and is not the target of this analysis while the order of convergence, $p$ is of interest and can be numerically approximated. The logarithm of the error is given by $\log\left(\|\mathbf{b}_{k}^{(N)}-\mathbf{b}^{\star}\|\right)=\log C-p\log N$.
@@
For example, consider the case where we double $N$ and take the log-ratio of the errors and take the difference of the logarithms: $\log_{2}\left(\|\mathbf{b}_{k}^{(N)}-\mathbf{b}^{\star}\|\right)-\log_{2}\left(\|\mathbf{b}_{k}^{(2N)}-\mathbf{b}^{\star}\|\right)=p.$ Running the algorithm multiple times on randomised datasets gives us an average estimate of $\|\mathbf{b}_{k}^{(N)}-\mathbf{b}^{\star}\|$ for a particular value of $N$. We also compute the $L^{2}$-error between the FEM interpolated functions $\|f_{k}-f^{\star}\|$, as a comparison. The order of convergence for two examples (1D Poisson on unit interval, 2D Poisson on disc) can be seen in Figure 1 and are estimated to be $(0.502,0.506)$ are close to $1/2$, which is the theoretical order of convergence derived in Theorem 1. We should also see the particle estimate of the posterior converge to the true posterior, $p(\mathbf{u}|\mathbf{y},\mathbf{b}^{\star})$; to assess this we compute the variance of the empirical estimate and compare this against the true posterior variance for increasing numbers of particles.
Figure 2: (left) True posterior $p(\mathbf{u}|\mathbf{y},\mathbf{b}^{\star})=\mathcal{N}(\mathbf{m}_{\mathbf{u}_{\mathbf{y}}},\mathbf{C}_{\mathbf{u}|\mathbf{y}})$, compared to particle approximated posterior $p(\mathbf{u}|\mathbf{y},\mathbf{b}_{k})=\frac{1}{N}\sum_{n=1}^{N}\delta_{\mathbf{u}_{k}^{(n)}}$, for $N=256$. (right) Variance of true posterior vs approximations for increasing number of particles.
@@ 6.2 Forcing estimation for Nonlinear Poisson PDE
For the nonlinear Poisson PDE (29), we introduce a known nonlinearity with a solution-dependent diffusivity $q(x)=1+u(x)^{2}$:
@@
6.3 Estimating Diffusivity Parameter for Linear Poisson PDE
For the inverse problem of estimating the diffusivity, we use the example in section 4.1:
$-\nabla\cdot(e^{\theta}\nabla u)$ $=f+\epsilon,\quad x\in(0,1),$
$u(x)$ $=0,\quad x\in\{0,1\}.$ (47)
$\nabla_{\mathbf{u}}\Psi^{y}_{\theta}(\mathbf{u})=e^{\theta}\mathbf{A}^{\top}\mathbf{G}^{-1}\left(e^{\theta}\mathbf{A}\mathbf{u}-\mathbf{b}\right)+\mathbf{H}^{\top}\mathbf{R}^{-1}\left(\mathbf{H}\mathbf{u}-\mathbf{y}\right).$
$\left\langle\mathbf{u}-\mathbf{u}^{\prime},\nabla_{\mathbf{u}}\Psi^{y}_{\theta_{0}}(\mathbf{u})-\nabla_{\mathbf{u}}\Psi^{y}_{\theta_{0}}(\mathbf{u}^{\prime})\right\rangle$ $\geq\mu\|\mathbf{u}-\mathbf{u}^{\prime}\|^{2},\quad\mu=\lambda_{\mathrm{min}}\left(e^{2\theta_{0}}\mathbf{A}^{\top}\mathbf{G}^{-1}\mathbf{A}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right),$
and Lipschitz continuity
$\left\|\nabla_{\mathbf{u}}\Psi^{y}_{\theta_{0}}(\mathbf{u})-\nabla_{\mathbf{u}}\Psi^{y}_{\theta_{0}}(\mathbf{u}^{\prime})\right\|$ $\leq L\|\mathbf{u}-\mathbf{u}^{\prime}\|,\quad L=\lambda_{\mathrm{max}}\left(e^{2\theta_{0}}\mathbf{A}^{\top}\mathbf{G}^{-1}\mathbf{A}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right),$
we can run the diffusion on $\mathbf{u}$ with fixed $\theta_{0}$ which will exhibit exponential convergence. Initialising $\mathbf{u}$ variables at stationarity for $\theta_{0}$ and running the system jointly produces a stable behaviour experimentally, which might be an interesting starting point of an analysis for this case. The particles are initialised at zero and the diffusivity parameter at some value larger than the truth. With shorter warm-starts, the parameter estimate increases rapidly from the initial value away from the true value because the particles have not had sufficient time to move away from zero, hence the parameter updates believe the system is highly diffusive. The second longest warm-start provides the quickest converge to the true diffusivity, with the longest spending too long in the warm-start stage. Figure 5 shows the parameter traces for these different lengths of warm-start, alongside the prior over the parameter $\theta$.
@@ 7 Conclusions
In this paper we have posed the problem of estimating PDE parameters as a MMAP problem, constructing statFEM generative model, to which we applied IPLA. Nonasymptotic bounds on the MMAP error have been proved, and experimentally confirmed for both 1D and 2D geometries. The approach in [4] has been extended to a nonlinear statFEM setting, and applied to the nonlinear Poisson PDE. Different methods for constructing Gaussian approximations to the non-Gaussian statFEM prior were considered and tested, with a particular focus on first-order Taylor approach. The effects of a ’warm-start’ were explored for the case of solving the inverse problem of estimating diffusivity, showing that a warm-start can improve the convergence of the estimate. With nonasymptotic bounds derived, further research will consider scaling up this approach to work with larger systems, where solving the forward problem is prohibitively expensive — in this case, Langevin dynamics can still be used to sample from the target density without forward solves [4]. Future work can also benefit from extensions to non-log-concave cases arising in nonlinear PDEs using the techniques introduced in [43, 2].

### v2

@@ 5 Experimental Results
In this section, we provide comprehensive numerical evidence that the methods we presented above are effective in practice. We consider two examples, with different levels of complexity, to demonstrate the effectiveness of the IPLA algorithm within the setting of inverse problems. Specifically, we consider the following examples:
@@
  • Linear Poisson inverse problem: We consider the estimation of an unknown constant diffusivity for the linear Poisson problem, demonstrating the effect of a warm-start on the efficiency of estimating the parameter.
@@ 5.1 Forcing estimation for Linear Poisson PDE
Here we consider two example domains, the unit interval, and a 2D disc. In both cases we fix the diffusivity to a constant value of one, i.e. $\varkappa=1$. For the unit interval the linear Poisson PDE is given by
@@
with $f(\mathbf{x})=100\exp\left(-\frac{\|\mathbf{x}+\mathbf{m}\|^{2}_{2}}{2*0.2^{2}}\right)+100\exp\left(-\frac{\|\mathbf{x}-\mathbf{m}\|^{2}_{2}}{2*0.2^{2}}\right)$, $\mathbf{m}=[0.4,0.4]^{\top}$, and $k(\mathbf{x},\mathbf{x}^{\prime})=\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}_{2}}{2*0.1^{2}}\right)$. For inference $f\sim\mathcal{GP}(0,k_{f})$, $k_{f}(\mathbf{x},\mathbf{x}^{\prime})=100\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}_{2}}{2*0.2^{2}}\right)$. For the examples in this work we assume the GP hyperparameters, such as kernel length-scale and amplitude are known a priori. For the forcing prior hyperparameters, these would be set by the practitioner to reflect the a priori knowledge. The GP misspecification hyperparameters would ideally be estimated from data. In this study we do not include such estimation, which could be performed by e.g. extending the MMLE/MMAP to include the kernel hyperparameters, via grid search, or Bayesian optimisation methods; we leave this as a consideration for future work.
@@
The convergence of the estimated forcing to the MMAP estimator is computed by assuming a convergence of the form $\|\mathbf{b}^{(N)}_{k}-\mathbf{b}^{\star}\|_{2}=CN^{-p}$, where $\mathbf{b}^{(N)}_{k}$ is the parameter estimate after $k$ iterations of preconditioned IPLA (27)–(28) using $N$ particles. The rate of convergence, $C$ will depend on the problem and is not the target of this analysis while the order of convergence, $p$ is of interest and can be numerically approximated. The logarithm of the error is given by $\log\left(\|\mathbf{b}_{k}^{(N)}-\mathbf{b}^{\star}\|\right)=\log C-p\log N$.
@@
For example, consider the case where we double $N$ and take the log-ratio of the errors and take the difference of the logarithms: $\log_{2}\left(\|\mathbf{b}_{k}^{(N)}-\mathbf{b}^{\star}\|\right)-\log_{2}\left(\|\mathbf{b}_{k}^{(2N)}-\mathbf{b}^{\star}\|\right)=p.$ Running the algorithm multiple times on randomised datasets gives us an average estimate of $\|\mathbf{b}_{k}^{(N)}-\mathbf{b}^{\star}\|$ for a particular value of $N$. We also compute the $L^{2}$-error between the FEM interpolated functions $\|f_{k}-f^{\star}\|$, as a comparison. The order of convergence for two examples (1D Poisson on unit interval, 2D Poisson on disc) can be seen in Figure 1 and are estimated to be $(0.502,0.506)$ are close to $1/2$, which is the theoretical order of convergence derived in Theorem 1. We should also see the particle estimate of the posterior converge to the true posterior, $p(\mathbf{u}|\mathbf{y},\theta,\mathbf{b}^{\star})$; to assess this we compute the variance of the empirical estimate and compare this against the true posterior variance for increasing numbers of particles, which we visualise in Figure 2 for the 1D Poisson example (5.1).
Figure 2: (left) True posterior $p(\mathbf{u}|\mathbf{y},\theta,\mathbf{b}^{\star})=\mathcal{N}(\mathbf{m}_{\mathbf{u}|\mathbf{y}},\mathbf{C}_{\mathbf{u}|\mathbf{y}})$, compared to particle approximated posterior $p(\mathbf{u}|\mathbf{y},\mathbf{b}_{k})=\frac{1}{N}\sum_{n=1}^{N}\delta_{\mathbf{u}_{k}^{(n)}}$, for $N=256$. (right) Variance of true posterior vs approximations for increasing number of particles.
@@
5.2 Estimating Diffusivity Parameter for Linear Poisson PDE
For the inverse problem of estimating the diffusivity, we use the 1D Poisson equation:
$-\nabla\cdot(e^{\theta}\nabla u)$ $=f+\epsilon,\quad x\in(0,1),$
$u(x)$ $=0,\quad x\in\{0,1\}.$ (38)
$\nabla_{\mathbf{u}}\Psi^{y}_{\theta}(\mathbf{u},\mathbf{b})=\mathbf{A}_{\theta}^{\top}\mathbf{G}^{-1}\left(\mathbf{A}_{\theta}\mathbf{u}-\mathbf{b}\right)+\mathbf{H}^{\top}\mathbf{R}^{-1}\left(\mathbf{H}\mathbf{u}-\mathbf{y}\right).$
$\left\langle\mathbf{u}-\mathbf{u}^{\prime},\nabla_{\mathbf{u}}\Psi^{y}_{\theta_{0}}(\mathbf{u},\mathbf{b})-\nabla_{\mathbf{u}}\Psi^{y}_{\theta_{0}}(\mathbf{u}^{\prime},\mathbf{b})\right\rangle$ $\geq\mu\|\mathbf{u}-\mathbf{u}^{\prime}\|^{2},$
and Lipschitz continuity
$\left\|\nabla_{\mathbf{u}}\Psi^{y}_{\theta_{0}}(\mathbf{u},\mathbf{b})-\nabla_{\mathbf{u}}\Psi^{y}_{\theta_{0}}(\mathbf{u}^{\prime},\mathbf{b})\right\|$ $\leq L\|\mathbf{u}-\mathbf{u}^{\prime}\|,$
where
$\mu=\lambda_{\mathrm{min}}\left(\mathbf{A}_{\theta_{0}}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta_{0}}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right),\quad L=\lambda_{\mathrm{max}}\left(\mathbf{A}_{\theta_{0}}^{\top}\mathbf{G}^{-1}\mathbf{A}_{\theta_{0}}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right).$
We can run the diffusion on $\mathbf{u}$ with fixed $\theta_{0}$ which will exhibit exponential convergence. Initialising $\mathbf{u}$ variables at stationarity for $\theta_{0}$ and running the system jointly produces a stable behaviour experimentally, which might be an interesting starting point of an analysis for this case. We refer to the iteration of the particle update (30) while fixing $\theta_{0}$ constant a ‘warm-start’, and the number of iterations for which this is run the ‘warm-start length’. The particles are initialised at zero and the diffusivity parameter at some value larger than the truth. With shorter warm-starts, the parameter estimate increases rapidly from the initial value away from the true value because the particles have not had sufficient time to move away from zero, hence the parameter updates believe the system is highly diffusive. The second longest warm-start provides the quickest converge to the true diffusivity, with the longest spending too long in the warm-start stage. Figure 3 shows the parameter traces for these different lengths of warm-start, alongside the prior over the parameter $\theta$.
6 Nonlinear Extension
As we have seen in section 2.1, if the PDE is defined by a linear differential operator the prior distribution induced by the GP noise is also a Gaussian. This is not the case in general. A nonlinear differential operator will induce a (typically non-analytic) non-Gaussian prior over the FEM coefficients. In this section we derive the statFEM prior for a nonlinear PDE, before deriving the potential and approximations for a nonlinear Poisson PDE, before demonstrating the forcing estimation results.
6.1 Statistical Finite Elements for Nonlinear PDEs
A stochastic PDE with the nonlinear differential operator $\mathcal{F}$ and forcing $f$
$\mathcal{F}(u)=f+\epsilon,\quad\epsilon\sim\mathcal{GP}(0,k(\cdot,\cdot)),$ (39)
can be discretised via the statFEM method giving a system of nonlinear algebraic equations $\mathbf{F}(\mathbf{u})=\mathbf{b}+\mathbf{e},\quad\mathbf{e}\sim\mathcal{N}(\mathbf{0},\mathbf{G})$ (see Appendix D for derivation). The induced prior distribution can be found via a transformation of variables [7], from distribution $\mathbf{E}\sim f_{\mathbf{E}}(\mathbf{e})$ to $\mathbf{U}\sim f_{\mathbf{U}}(\mathbf{u})$ where $\mathbf{u}=g(\mathbf{e})$,
Using this transformation of variables with $f_{\mathbf{E}}=\mathcal{N}(\mathbf{0},\mathbf{G})$ and $g^{-1}(\mathbf{u})=\mathbf{F}(\mathbf{u})-\mathbf{b}$, we can see how the resulting density of the solution depends on the nonlinear differential operator and its Jacobian:
$p(\mathbf{u}|\mathbf{b})=\det|\mathbf{J}(\mathbf{u})|\frac{1}{Z}\exp\left\{-\frac{1}{2}\left(\mathbf{F}(\mathbf{u})-\mathbf{b}\right)^{\top}\mathbf{G}^{-1}\left(\mathbf{F}(\mathbf{u})-\mathbf{b}\right)\right\},$ (41)
where the constant $Z=(2\pi)^{-n_{u}/2}\det|\mathbf{G}|^{-1/2}$, and the Jacobian of the nonlinear transform $\mathbf{F}$ is denoted $\mathbf{J}(\mathbf{u})=\frac{\partial\mathbf{F}}{\partial\mathbf{u}}\big{|}_{\mathbf{u}}$. This prior, in general, is non-Gaussian due to the Jacobian determinant term.
@@ 6.2 Forcing Estimation for Nonlinear Poisson PDE
$-\nabla\cdot(q(u)\nabla u)=f+\epsilon.$ (42)
For full details see Appendix D. With a basis function expansion $u_{h}=\sum_{i=1}^{n_{u}}\hat{u}_{i}\phi_{i}$, this is rewritten in vector form
where we have $\mathbf{u}=\left[\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{n_{u}}\right]^{\top}$, $\left[\mathbf{F}(\mathbf{u})\right]_{i}=\left\langle q(u_{h})\nabla u_{h},\nabla\phi_{i}\right\rangle$, $\left[\mathbf{b}\right]_{i}=\langle f,\phi_{i}\rangle$, and $\left[\mathbf{G}\right]_{ij}=\langle\phi_{i},\langle k(\cdot,\cdot),\phi_{j}\rangle\rangle$. The prior potential we define as $\Phi(\mathbf{u},\mathbf{b})\coloneqq-\log p(\mathbf{u}|\mathbf{b})$, and posterior potential $\Phi^{y}(\mathbf{u},\mathbf{b})\coloneqq\Phi(\mathbf{u},\mathbf{b})-\log p(\mathbf{y}|\mathbf{u})$ using the same observation model as (2). Finally, in the same manner as the linear Poisson case we use a GP prior for the forcing function, and define the joint potential as $\Psi^{y}(\mathbf{u},\mathbf{b})\coloneqq\Phi^{y}(\mathbf{u},\mathbf{b})-\log p(\mathbf{b})$.
In this nonlinear case the computation of the gradient of the prior potential with respect to $\mathbf{u}$, $\nabla_{\mathbf{u}}\Phi(\mathbf{u},\mathbf{b})$ is intractable due to the log-determinant Jacobian term in the score of the nonlinear statFEM prior, so we turn to approximations of this distribution to make progress. Approximating this distribution with a Gaussian alleviates the problem of intractability, providing a linear score computed using the approximate mean and covariance. We present here the algorithm for sampling based on a Gaussian approximation of the nonlinear statFEM prior $p(\mathbf{u}|\mathbf{b}_{k})\approx\mathcal{N}(\mathbf{m}_{k},\mathbf{C}_{k})$:
$\mathbf{b}_{k+1}$ $=(1-\gamma)\mathbf{b}_{k}+\gamma\mathbf{P}_{\mathbf{b}}\left(\mathbf{G}^{-1}\left(\frac{1}{N}\sum_{n=1}^{N}\mathbf{F}(\mathbf{u}_{k}^{(n)})\right)+\bm{\Sigma}^{-1}\bm{\mu}\right)+\sqrt{\frac{2\gamma}{N}}\mathbf{P}_{\mathbf{b}}^{1/2}\zeta_{k+1}^{(0)}$ (45)
$\mathbf{u}_{k+1}^{(n)}$ $=(1-\gamma)\mathbf{u}_{k}^{(n)}+\gamma\mathbf{P}_{\mathbf{u}}\left(\mathbf{C}_{k}^{-1}\mathbf{m}_{k}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{y}\right)+\sqrt{2\gamma}\mathbf{P}_{\mathbf{u}}^{1/2}\zeta_{k+1}^{(n)},\quad\forall n\in\left\{1,\dots,N\right\}$ (46)
$\text{with}\quad\mathbf{P}_{\mathbf{b}}=\left[\mathbf{G}^{-1}+\bm{\Sigma}^{-1}\right]^{-1},\quad\mathbf{P}_{\mathbf{u}}=\left[\mathbf{C}_{k}^{-1}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right]^{-1}.$ (47)
Note the gradient with respect to the forcing can be computed exactly because the potential is quadratic in $\mathbf{b}$, i.e. $\nabla_{\mathbf{b}}\Psi^{y}(\mathbf{u},\mathbf{b})=\mathbf{G}^{-1}\left(\mathbf{F}(\mathbf{u})-\mathbf{b}\right)$. Simple approaches to approximating the distribution of nonlinearly transformed Gaussian random variables [29] include a first-order Taylor approximation, the unscented transform, and an empirical approximation to the first two moments with a sample based approach (See Table 3). The update steps (45)–(46) require the computation of the preconditioner at each iteration as the approximation to $p(\mathbf{u}|\mathbf{b}_{k})$ is updated. This is a computationally expensive step, which requires the matrix inverse $\mathbf{C}_{k}^{-1}$ requiring order $n_{u}^{3}$ operations, followed by the inverse $\left[\mathbf{C}_{k}^{-1}+\mathbf{H}^{\top}\mathbf{R}^{-1}\mathbf{H}\right]^{-1}$, again order $n_{u}^{3}$. We note that for first-order Taylor, we can compute the inverse of covariance matrix efficiently from the Jacobian by noting $\mathbf{C}_{k}^{-1}=\mathbf{J}(\mathbf{m}_{k})^{\top}\mathbf{G}^{-1}\mathbf{J}(\mathbf{m}_{k})$ which is only order $n_{u}^{2}$. For larger-scale problems when this may become infeasible to recompute the preconditioner at each iteration other methods can be used for preconditioning, since for the first-order Taylor method, the preconditioning computation is order $n_{u}^{3}$ but the approximation is order $n_{u}^{2}$, hence fixing this preconditioner for some iterations may have significant impacts on efficiency. For example, a fixed preconditioned could be used, computed at the prior mean, or the preconditioner could be recomputed after a fixed number of steps. These different preconditioning strategies are of interest in future work.
We remark that all these methods require us to solve the nonlinear system, and the first-order Taylor method requires computation of the Jacobian. This procludes the use of the first-order Taylor method for non-differentiable nonlinearities, where we can only evaluate $\mathbf{F}$ and not its Jacobian $\mathbf{J}$; for this case, fixed-point methods could be used to solve $\mathbf{F}(\mathbf{u})=\mathbf{b}$ and the unscented transform or Monte Carlo methods applied for approximating the covariance. For this example, since we have access to the Jacobian, we apply Newton’s method for solving the system.
Table 3: Gaussian approximations to nonlinearly transformed Gaussian random variables: first-order Taylor (FOT), unscented transform (UT) and Monte Carlo (MC). FOT requires one solve followed by computation of the Jacobian, $\mathbf{J}(\hat{\mathbf{m}})=\frac{\partial\mathbf{F}}{\partial\mathbf{u}}|_{\hat{\mathbf{m}}}$. UT requires $12n_{u}+$ 1 solves, details in Appendix E.2. MC requires $M$ solves, where $\tilde{\mathbf{b}}^{(j)}\sim\mathcal{N}(\mathbf{b},\mathbf{G})$.
Solve Mean: $\mathbf{m}$ Covariance: $\mathbf{C}$
Details of each of these approximations are given in Appendix E.2.
6.3 Nonlinear Poisson Experimental Results
This example introduces a nonlinearity into the Poisson diffusivity, where we can compare different nonlinear approximation methods. Here, we compare the first-order Taylor, unscented transform and Monte Carlo approximations for approximating nonlinear transforms of Gaussian random variables. For the nonlinear Poisson PDE (42), we introduce a known nonlinearity with a solution-dependent diffusivity $q(x)=1+u(x)^{2}$:
@@ 7 Conclusions
In this paper we have posed the problem of estimating PDE parameters as a MMAP problem, constructing statFEM generative model, to which we applied IPLA. Nonasymptotic bounds on the MMAP error have been proved, and experimentally confirmed for both 1D and 2D geometries. The approach in [4] has been extended to a nonlinear statFEM setting, and applied to the nonlinear Poisson PDE. Different methods for constructing Gaussian approximations to the non-Gaussian statFEM prior were considered and tested, with a particular focus on first-order Taylor approach. The effects of a ’warm-start’ were explored for the case of solving the inverse problem of estimating diffusivity, showing that a warm-start can improve the convergence of the estimate. With nonasymptotic bounds derived, further research will consider scaling up this approach to work with larger systems, where solving the forward problem is prohibitively expensive — in this case, Langevin dynamics can still be used to sample from the target density without forward solves [4]. Future work can also benefit from extensions to non-log-concave cases arising in nonlinear PDEs using the techniques introduced in [51, 2]. Recent developments in particle-based schemes for accelerating the gradient descent are of interest in future work, such as the underdamped Langevin extension to IPLA [38], the momentum-based approach of [35], and the Stein variational gradient descent method of [43].